OpenStudy (anonymous):

Please factor the following polynomial completely 208p^2-325q^2

OpenStudy (anonymous):

helps to know that both numbers have a common factor of 13

OpenStudy (anonymous):

so start with \[13(16p^2-25q^2)\] and once again factor as the difference of two square \[a^2-b^2=(a+b)(a-b)\] this time with \[a=6p, b=5q\]

OpenStudy (anonymous):

sorry \[a=4p, b=5q\]

OpenStudy (callisto):

hmm, continue : \[13(16^{2}-25q ^{2}) =13(4p-5q)(4p+5q)\]